OpenStudy (anonymous):

how do you get an answer to Fg=m1m2/d2

OpenStudy (anonymous):

Do you mean \[F_g=G\frac{m_1m_2}{d^2}\]Is there a specific problem you're trying to apply it to?

OpenStudy (anonymous):

yes thats it and i have a test on it and i dont have a clue on how to to do it

OpenStudy (anonymous):

what was the test question?

OpenStudy (anonymous):

no i have a test tomorrow

OpenStudy (anonymous):

I apologize if I'm starting too far back...do you know what the variables represent?

OpenStudy (anonymous):

are you reviewing from a practice test or something?

OpenStudy (anonymous):

yes im reviewing and i know what the variables mean

OpenStudy (anonymous):

Ok, I didn't want to be condescending, but I also don't want to assume anything. Give me a minute and I'll see if I can find a representative problem for an example.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):



OpenStudy (anonymous):

two masses with their centers a distance apart of 10,000m m2 is 20,000kg m1 is 30,000kg what's the gravitational force either mass experiences?

OpenStudy (anonymous):

fg=30000*20000 ------------- 10000 * 10000 fg=6

OpenStudy (anonymous):

not bad, you forgot the constant though...

OpenStudy (anonymous):

G*6

OpenStudy (anonymous):

Almost, you need the gravitational constant.

OpenStudy (anonymous):

Sheesh. Like racing a machine. :P

OpenStudy (anonymous):

look at the equation he showed you... see the giant G ?

OpenStudy (anonymous):

it's a constant, check your text for the value.

OpenStudy (anonymous):

6*9.8=58.8

OpenStudy (anonymous):

not that constant...

OpenStudy (anonymous):

capital G

OpenStudy (anonymous):

6.67300 × 10-11 m3 kg-1 s-2

OpenStudy (anonymous):

ok thanks i get it